# DialSim: A Real-Time Simulator for Evaluating Long-Term Multi-Party Dialogue Understanding of Conversation Systems

### v5

@@ 6 Conclusion
In this paper, we introduce DialSim, a simulator designed for evaluating the ability of conversation systems to understand long-term, multi-party dialogues in real-time settings. To run DialSim, we first constructed LongDialQA, a dataset based on dialogues from well-known TV show scripts. LongDialQA also includes questions derived from fan quizzes and a temporal knowledge graph, enabling a comprehensive assessment of conversation systems. Using DialSim, we evaluated the latest conversation systems and uncovered significant limitations in their ability to effectively handle complex, multi-party, long-term dialogues in real-time scenarios.
Limitations
Despite its strengths, our simulator has two main limitations. First, while the questions and answers are logically paired for accurate evaluation, the random selection of questions could introduce a bit of awkwardness during conversations. Second, while we considered incorporating industry-specific dialogues such as chat logs from customer service or retail, where conversation systems could be used for business purposes, these dialogue datasets are usually proprietary and not publicly accessible. In future developments, we will focus on enhancing the natural flow of interactions and creating simulators that are applicable to real-world industries.
@@ Appendix F Prompt for Response Generation
The prompt given to the conversation system to answer questions using dialogue history is shown in Table 9. An example where the placeholders from Table 9 are filled with actual values can be found in Table 10.
@@ Appendix G Experimental Setting
G.1 Time Limit
In DialSim, the time limit is a controllable parameter, giving developers the flexibility to conduct experiments with any chosen time constraint, or even without one. When a time limit is set, the experimental environment can impact performance. Consequently, depending on the environment in which the conversation system is deployed, this could serve as a criterion for selecting the system with relatively better performance. It is important to note that the primary objective of DialSim is not to evaluate the inference speed of LLMs, but rather to assess the end-to-end performance of conversation systems, where techniques like model sharding and tensor parallelism can be a part of the conversation system to decrease the response latency if needed.
To control the environmental factors that could affect time, we conducted all experiments under the same conditions as described in Appendix G.1.1. The rationale for setting a 6-second time limit in our experiments is detailed in Appendix G.1.2, and an analysis of the Internet speed for API-based models can be found in Appendix G.1.3.
G.1.1 Environment Control
Our simulator operates in real-time, requiring precise control of the experimental environment. Therefore, we conducted all experiments using the same hardware: NVIDIA RTX A6000 GPUs and an AMD EPYC 7702 64-Core Processor. To maintain consistent CPU performance, we allocated 10 cores for each experiment and ensured that no other processes were running simultaneously.
G.1.2 Average Time Interval Between Utterances
Each episode includes around 240 utterances and lasts about 18 minutes without commercial breaks. This means each utterance should occur roughly every 4.5 seconds. However, because the experiments used the A6000, which is slower than the latest hardware like the A100 or H100, we extended the interval to 6 seconds.
To account for this, we set a 6-second window as the response time limit for systems and conducted experiments to determine whether current models could meet this criterion. It is important to emphasize that the primary goal of these experiments was not to evaluate the absolute performance of the models but to showcase the range of analyses possible under time limits.
G.1.3 Internet Speed
The performance of API-based models can be affected by internet speed. To analyze this, we conducted a comparative analysis of the response times between API-based models and open-source models. In our analysis of systems using OpenAI Embedding-Session Sum., we found that the API-based systems achieved average response times of 1.50 seconds for GPT-4o-mini, 1.73 seconds for GPT-3.5 and 2.69 seconds for Gemini 1.0 pro. In comparison, systems using open-source models showed average response times ranging from 2.06 seconds (Gemma 2B) to 7.15 seconds (Tulu2 70B). These results suggest that, even when accounting for both internet communication and model inference, remote API-based models are generally faster than open-source alternatives. This indicates that internet latency has a minimal impact on our evaluation.
G.2 Question Format
LongDialQA is a dataset that includes pairs of questions, answers, and choices. The questions are available in three formats: template-based multiple-choice, natural language multiple-choice, and open-ended. Users can choose any of these formats to evaluate the system’s performance.
@@ G.2 Question Format
Additionally, we offer the option to ask questions in an open-ended format (e.g., “Who was going out with Paul in September 1994?”) without providing answer choices. This approach allows us to evaluate the system’s ability to generate open-ended responses. The open-ended format is particularly useful for fan quiz-based questions, where some answers may require longer responses (e.g., Question: “Why did Monica and Chandler say they were late getting to the hospital?” Correct answer: “Monica went back for her jacket”).
@@
G.2.1 Choices in Multiple-Choice Questions
@@
G.3 Number of Retrieved Dialogue History
By default, systems retrieved up to 20 utterances, 10 entire sessions, and 15 session summaries, depending on the storing method, though some LLMs with shorter context lengths retrieved fewer histories accordingly.
@@ Appendix H Experimental Results for The Big Bang Theory and The Office
Appendix I Experimental Results in an Unlimited Time Setting
The experimental results for the unlimited time setting are presented in Table 13.
Appendix J Experimental Results for Different Time Limits
The experimental results for different time limits are shown in Figure 6 and Figure 7. Figure 6 illustrates the performance over different time limits in the BM25-Session Entire setting, while Figure 7 displays the performance in the Oracle setting. Due to the high costs, time-based experiments with GPT-4o, Gemini-1.5 Pro, and Claude-3 Opus were conducted exclusively in the Oracle setting. One key observation from the results is the performance of GPT-3.5, GPT-4o-mini, and GPT-4o. These models demonstrated consistent performance with quick inference times, handling up to a 3-second limit in the BM25-Session Entire setting and up to a 1-second limit in the Oracle setting. Consequently, these models are optimal for tasks requiring real-time communication without delays.
Appendix K Experimental Results in the Oracle Setting
Figure 8 shows the performance comparison between the BM25-Session Entire setting and the Oracle setting. These experiments were conducted without a time limit. Llama3.1-70B achieved the highest performance with a score of 69.86% in the Oracle setting.
Appendix L Experimental Results on Adversarial Test
In the adversarial test, we altered the characters’ names and ran experiments under different conditions. Table 14 displays the results when characters’ names were mixed with a 6-second time limit, while Table 15 shows the results without a time limit. Table 16 presents the results of changing characters’ names to new ones with a 6-second time limit, while Table 17 shows the results without a time limit.
Appendix M Annotator Instructions
@@ Appendix I Experimental Results in an Unlimited Time Setting
Question Type Relation Template Question Example
[…]
@@
[…]
@@
roommate, wife, husband, girlfriend, boyfriend, client, neighbor, boss, subordinate, fiance, fiancee
@@
mother, father, son, daughter, sister, brother
@@
Table 7: Templates for two-hop questions.
@@
By the way, Rachel had a roommate on October 28, 1994. Who dated the roommate in September 1994? Default
@@
Table 8: Examples of the results of character style transfer.
@@
You are $<$Chatbot$>$, a long-term conversational agent capable of interacting with multiple users.
@@
2. Your responses should solely rely on the retrieved dialogue history. If the information in the dialogue
history is insufficient to answer the question, you must choose (E).
3. This question is being asked in the context of $<$Date$>$.
@@
$<$Dialog_History$>$
[Question] $<$Question$>$
@@
Table 9: In the $<$Chatbot$>$ placeholder, the name of the main character (i.e., Ross, Sheldon, Michael) for each TV show is inserted. In the $<$Date$>$ placeholder, the date of the session in which the question is being asked is inserted. In the $<$Dialog_History$>$ placeholder, the dialogue history that the agent will use is inserted. In the $<$Question$>$ placeholder, the question that the agent should answer along with five choices is inserted.
@@
2. Your responses should solely rely on the retrieved dialogue history. If the information in the dialogue
history is insufficient to answer the question, you must choose (E).
@@
Monica: Okay, umm-umm, I’ll just–I’ll be right back, I just gotta go ah, go ah…
@@
Table 10: An actual example of the prompt for response generation.
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API GPT-4o-mini 22.68 (2.12) 19.77 (2.02) 36.63 (1.82) 30.10 (1.44) 29.54 (0.71) 32.34 (0.58) 35.72 (0.81) GPT-3.5 32.49 (1.72) 25.32 (1.20) 35.59 (2.12) 33.86 (1.09) 27.81 (0.40) 32.97 (0.86) 37.02 (1.13) Gemini 1.0 pro 3.49 (0.69) 25.87 (1.23) 30.72 (0.18) 38.16 (1.25) 37.42 (0.68) 32.09 (0.44) 36.30 (0.32) Open Tülu 2-70B 0.62 (0.13) 21.08 (0.70) 18.95 (1.07) 22.36 (0.65) 34.64 (0.69) 9.08 (1.00) 20.22 (1.48) Tülu 2-7B 0.53 (0.18) 15.58 (1.34) 22.26 (0.53) 29.99 (0.57) 16.84 (2.13) 21.48 (0.77) 28.69 (1.15) Llama3.1-70B 0.25 (0.07) 21.55 (0.93) 0.15 (0.12) 1.26 (0.31) 34.21 (1.59) 3.89 (1.05) 14.89 (1.74) Llama3.1-8B 21.30 (1.68) 12.80 (1.06) 25.50 (0.16) 18.56 (0.99) 23.10 (2.69) 25.48 (3.65) 20.75 (1.58) Mixtral-8x7B 1.95 (0.34) 15.91 (0.71) 34.52 (1.12) 16.83 (1.60) 17.45 (0.49) 34.98 (0.99) 13.83 (2.18) Mistral-7B 3.11 (0.21) 24.69 (1.82) 34.26 (0.60) 32.17 (1.39) 30.23 (0.62) 33.36 (0.56) 29.19 (1.54) Gemma-7B 16.40 (0.74) 21.40 (2.33) 19.74 (2.45) 16.67 (0.40) 24.50 (1.87) 20.22 (1.39) 16.12 (0.52) Gemma-2B 1.56 (0.06) 28.94 (0.35) 26.12 (2.22) 33.47 (1.41) 27.92 (0.68) 29.40 (1.79) 34.86 (3.20)
$\dagger$: Both GPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
Table 11: The performances of the agents on The Big Bang Theory dialogue in DialSim (time limit = 6 seconds). We conducted experiments three times and reported the accuracies and the standard deviations.
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API GPT-4o-mini 28.48 (1.01) 29.44 (0.62) 43.16 (1.37) 35.92 (2.50) 37.81 (0.30) 40.91 (0.37) 42.83 (1.12) GPT-3.5 36.54 (0.32) 36.63 (0.57) 45.33 (1.00) 40.93 (0.13) 42.49 (1.24) 43.04 (0.82) 45.18 (0.56) Gemini 1.0 pro 2.42 (0.18) 35.11 (0.50) 48.90 (1.57) 40.91 (0.75) 44.72 (0.19) 46.63 (0.89) 45.82 (0.97) Open Tülu 2-70B 0.46 (0.09) 22.33 (1.00) 35.52 (0.89) 23.49 (1.16) 38.61 (1.02) 43.49 (1.27) 23.54 (0.52) Tülu 2-7B 0.32 (0.04) 25.86 (0.54) 27.95 (1.03) 36.60 (2.11) 22.13 (0.33) 29.50 (0.56) 35.51 (1.40) Llama3.1-70B 0.19 (0.07) 29.21 (0.56) 13.31 (0.94) 21.32 (5.22) 47.41 (0.93) 47.07 (1.32) 19.46 (1.73) Llama3.1-8B 21.87 (0.60) 22.03 (0.32) 37.94 (1.28) 29.16 (1.80) 27.76 (3.52) 37.67 (1.70) 26.67 (0.83) Mixtral-8x7B 1.53 (0.41) 19.63 (0.79) 34.35 (1.19) 16.07 (0.56) 20.02 (0.44) 30.44 (1.69) 12.43 (1.04) Mistral-7B 2.55 (0.09) 30.65 (0.45) 41.16 (1.26) 35.67 (1.68) 36.92 (2.13) 42.71 (1.24) 37.65 (2.42) Gemma-7B 17.81 (0.86) 21.58 (0.61) 25.62 (0.02) 12.20 (0.57) 24.88 (0.93) 24.38 (0.52) 15.70 (0.43) Gemma-2B 0.83 (0.16) 29.71 (0.69) 28.11 (1.14) 34.63 (0.94) 31.54 (0.65) 30.31 (0.16) 33.37 (0.27)
$\dagger$: Both GPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
Table 12: The performances of the agents on The Office dialogue in DialSim (time limit = 6 seconds). We conducted experiments three times and reported the accuracies and the standard deviations.
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API GPT-4o-mini 38.91 (0.99) 34.44 (0.52) 49.21 (0.12) 42.23 (1.57) 38.91 (0.74) 43.64 (0.42) 42.40 (0.99) GPT-3.5 31.81 (1.33) 26.91 (2.30) 39.45 (1.40) 32.77 (1.31) 32.41 (0.96) 35.78 (0.74) 35.98 (1.75) Gemini 1.0 pro 28.36 (0.97) 28.10 (1.08) 39.90 (1.08) 34.11 (1.64) 34.26 (2.91) 30.93 (2.17) 33.96 (2.11) Open Tülu 2-70B 3.31 (0.32) 29.87 (0.65) 35.87 (2.56) 34.72 (1.63) 37.07 (0.72) 33.63 (1.32) 38.62 (1.94) Tülu 2-7B 1.57 (0.12) 28.93 (2.81) 28.72 (1.80) 30.86 (2.29) 34.55 (0.47) 31.04 (0.96) 32.12 (0.75) Llama3.1-70B 36.36 (0.68) 31.84 (1.29) 43.17 (0.99) 43.81 (0.94) 39.85 (2.08) 43.17 (0.68) 48.49 (0.97) Llama3.1-8B 28.78 (0.34) 29.89 (1.56) 34.70 (1.75) 33.93 (1.76) 31.63 (2.17) 32.91 (0.51) 35.59 (1.09) Mixtral-8x7B 42.19 (1.76) 31.84 (0.78) 46.47 (1.75) 32.31 (1.09) 35.51 (0.19) 41.24 (2.90) 34.18 (0.96) Mistral-7B 32.93 (0.59) 28.20 (1.17) 35.09 (1.76) 30.16 (1.82) 30.12 (1.45) 31.00 (1.93) 30.80 (1.75) Gemma-7B 18.78 (0.87) 22.26 (1.52) 23.62 (2.09) 19.83 (1.74) 25.07 (0.49) 22.48 (0.25) 20.08 (0.76) Gemma-2B 1.16 (0.26) 25.03 (1.54) 24.64 (1.31) 24.84 (2.05) 28.06 (1.38) 24.56 (2.60) 28.28 (1.94)
$\dagger$: Both GPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
@@
Table 13: The performance of the agents on Friends dialogue in DialSim (without time limit). We conducted experiments three times and reported the accuracy and standard deviations.
Figure 6: The experimental results for different time limits in the BM25-Session Entire setting.
Figure 7: The experimental results for different time limits in the Oracle setting.
@@
Figure 8: The performance comparison between the BM25-Session Entire setting and the Oracle setting.
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API GPT-4o-mini 31.37 (1.57) 27.54 (1.85) 40.36 (0.73) 29.71 (0.84) 33.63 (0.30) 35.93 (1.56) 37.30 (1.71) Open Tülu 2-70B 0.31 (0.13) 17.03 (0.94) 15.20 (0.87) 18.45 (1.04) 26.89 (0.54) 6.92 (0.47) 12.86 (1.23) Tülu 2-7B 0.73 (0.29) 12.50 (1.73) 17.58 (1.14) 24.21 (1.09) 10.20 (0.21) 14.26 (0.92) 21.03 (0.58) Llama3.1-70B 0.51 (0.00) 27.84 (1.89) 0.60 (0.06) 13.84 (2.31) 35.67 (1.89) 1.53 (028) 20.90 (0.37) Llama3.1-8B 25.84 (1.16) 25.24 (0.30) 28.86 (1.01) 24.56 (0.99) 28.86 (1.10) 32.35 (1.51) 24.05 (1.36) Mixtral-8x7B 1.77 (0.19) 14.03 (0.12) 21.11 (1.07) 15.50 (0.68) 13.14 (0.83) 18.03 (0.44) 18.47 (0.55) Mistral-7B 2.34 (0.17) 22.29 (1.43) 27.08 (0.99) 24.15 (1.76) 25.17 (1.74) 26.76 (2.64) 23.81 (2.53) Gemma-7B 18.87 (1.43) 22.85 (0.81) 22.96 (1.34) 17.95 (0.62) 25.46 (2.08) 21.53 (1.00) 17.66 (1.31) Gemma-2B 0.78 (0.22) 22.99 (0.66) 25.48 (1.54) 25.86 (2.48) 25.08 (1.34) 25.21 (0.22) 26.14 (1.71)
$\dagger$: Both GPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.

### v6

@@ 6 Conclusion
In this paper, we introduce DialSim, a simulator designed for evaluating the ability of conversational agents to understand long-term, multi-party dialogues. To run DialSim, we first constructed LongDialQA, a dataset based on dialogues from well-known TV show scripts. LongDialQA also includes questions derived from fan quizzes and a temporal knowledge graph, enabling a comprehensive assessment of the agents. Using DialSim, we evaluated the latest conversational agents and uncovered significant limitations in handling complex, multi-party, long-term dialogues. We hope our work paves the way for more rigorous and realistic evaluation standards in conversational AI.
@@ Appendix F Prompt for Response Generation
The prompt given to the conversational agent to answer questions using dialogue history is shown in Table 9. An example where the placeholders from Table 9 are filled with actual values can be found in Table 10.
@@ Appendix G Experimental Setting
G.1 Question Format
LongDialQAis a dataset that includes pairs of questions, answers, and choices. The questions are available in three formats: template-based multiple-choice, natural language multiple-choice, and open-ended. Users can choose any of these formats to evaluate the agent’s performance.
@@ G.1 Question Format
Additionally, we offer the option to ask questions in an open-ended format (e.g., “Who was going out with Paul in September 1994?”) without providing answer choices. This approach allows us to evaluate the agent’s ability to generate open-ended responses. The open-ended format is particularly useful for fan quiz-based questions, where some answers may require longer responses (e.g., Question: “Why did Monica and Chandler say they were late getting to the hospital?” Correct answer: “Monica went back for her jacket”).
@@
G.1.1 Choices in Multiple-Choice Questions
@@
G.2 Number of Retrieved Dialogue History
By default, agents retrieved up to 20 utterances, 10 entire sessions, and 15 session summaries, depending on the storing method, though some LLMs with shorter context lengths retrieved fewer histories accordingly.
@@ Appendix H Experimental Results for The Big Bang Theory and The Office
Appendix I Experimental Results in the Oracle Setting
Figure 7 shows the performance comparison between the BM25-Session Entire setting and the Oracle setting. Gemini-2.5 Flash achieved the highest performance with a score of 75.10% in the Oracle setting.
Appendix J Experimental Results on Adversarial Test
In the adversarial test, we altered the characters’ names and ran experiments under different conditions. Table 13 presents the results of original character names. Table 14 displays the results when characters’ names were mixed.
Appendix K Annotator Instructions
@@ Appendix I Experimental Results in the Oracle Setting
Without Time alumni Who is $\{\text{sub}\}$’s alumni? Who is Lincoln High School’s alumni?
boss Who is $\{\text{sub}\}$’s boss? Who is Chandler’s boss?
subordinate Who is $\{\text{sub}\}$’s subordinate? Who is Chandler’s subordinate?
client Who is $\{\text{sub}\}$’s client? Who is Chandler’s client?
neighbor Who is $\{\text{sub}\}$’s neighbor? Who is Chandler’s neighbor?
roommate Who is $\{\text{sub}\}$’s roommate? Who is Chandler’s roommate?
ex-roommate Who is $\{\text{sub}\}$’s ex-roommate? Who is Chandler’s ex-roommate?
fiance Who is $\{\text{sub}\}$’s fiance? Who is Rachel’s fiance?
fiancee Who is $\{\text{sub}\}$’s fiancee? Who is Ross’s fiancee?
ex-fiance Who is $\{\text{sub}\}$’s ex-fiance? Who is Rachel’s ex-fiance?
ex-fiancee Who is $\{\text{sub}\}$’s ex-fiancee? Who is Ross’s ex-fiancee?
pet Who is $\{\text{sub}\}$’s pet? Who is Ross’s pet?
dating with Who is dating $\{\text{sub}\}$? Who is dating Ross?
job What is $\{\text{sub}\}$’s job? What is Ross’s job?
place of work Where does $\{\text{sub}\}$ work? Where does Ross work?
age How old is $\{\text{sub}\}$? How old is Ross?
major What is $\{\text{sub}\}$’s major? What is Ross’s major?
mother Who is $\{\text{sub}\}$’s mother? Who is Ross’s mother?
father Who is $\{\text{sub}\}$’s father? Who is Ross’s father?
place of birth Where was $\{\text{sub}\}$ born? Where was Ben born?
hometown Where is $\{\text{sub}\}$’s hometown? Where is Monica’s hometown?
date of birth When was $\{\text{sub}\}$ born? When was Ben born?
husband Who is $\{\text{sub}\}$’s husband? Who is Emily’s husband?
wife Who is $\{\text{sub}\}$’s wife? Who is Ross’s wife?
girlfriend Who is $\{\text{sub}\}$’s girlfriend? Who is Joey’s girlfriend?
boyfriend Who is $\{\text{sub}\}$’s boyfriend? Who is Monica’s boyfriend?
ex-husband Who is $\{\text{sub}\}$’s ex-husband? Who is Carol’s ex-husband?
ex-wife Who is $\{\text{sub}\}$’s ex-wife? Who is Ross’s ex-wife?
ex-girlfriend Who is $\{\text{sub}\}$’s ex-girlfriend? Who is Ross’s ex-girlfriend?
ex-boyfriend Who is $\{\text{sub}\}$’s ex-boyfriend? Who is Rachel’s ex-boyfriend?
brother Who is $\{\text{sub}\}$’s brother? Who is Monica’s brother?
sister Who is $\{\text{sub}\}$’s sister? Who is Ross’s sister?
@@
Question Type Relation Template Question Example
With Time boss Who was $\{\text{sub}\}$’s boss on $\{\text{time}\}$? Who was Chandler’s boss on September 26th, 1994?
client Who was $\{\text{sub}\}$’s client on $\{\text{time}\}$? Who was Chandler’s client on September 26th, 1994?
neighbor Who was $\{\text{sub}\}$’s neighbor on $\{\text{time}\}$? Who was Chandler’s neighbor on September 26th, 1994?
roommate Who was $\{\text{sub}\}$’s roommate on $\{\text{time}\}$? Who was Chandler’s roommate on September 26th, 1994?
fiance Who was $\{\text{sub}\}$’s fiance on $\{\text{time}\}$? Who was Rachel’s fiance on September 26th, 1994?
fiancee Who was $\{\text{sub}\}$’s fiancee on $\{\text{time}\}$? Who was Ross’s fiancee on September 26th, 1994?
pet Who was $\{\text{sub}\}$’s pet on $\{\text{time}\}$? Who was Ross’s pet on September 26th, 1994?
dating with Who dated $\{\text{sub}\}$ on $\{\text{time}\}$? Who dated Ross on September 26th, 1994?
job What was $\{\text{sub}\}$’s job on $\{\text{time}\}$? What was Monica’s job on September 26th, 1994?
place of work Where did $\{\text{sub}\}$ work on $\{\text{time}\}$? Where did Monica work on September 26th, 1994?
age How old was $\{\text{sub}\}$ on $\{\text{time}\}$? How old was Monica on September 26th, 1994?
major What was $\{\text{sub}\}$’s major on $\{\text{time}\}$? What was Ross’s major on September 26th, 1994?
husband Who was $\{\text{sub}\}$’s husband on $\{\text{time}\}$? Who was Emily’s husband on September 26th, 1994?
wife Who was $\{\text{sub}\}$’s wife on $\{\text{time}\}$? Who was Ross’s wife on September 26th, 1994?
girlfriend Who was $\{\text{sub}\}$’s girlfriend on $\{\text{time}\}$? Who was Ross’s girlfriend on September 26th, 1994?
boyfriend Who was $\{\text{sub}\}$’s boyfriend on $\{\text{time}\}$? Who was Rachel’s boyfriend on September 26th, 1994?
Table 7: Templates for two-hop questions.
@@
roommate, wife, husband,
girlfriend, boyfriend, client,
neighbor, boss, subordinate,
fiance, fiancee
@@
mother, father, son,
daughter, sister, brother
@@
Table 8: Examples of the results of character style transfer.
@@
By the way, Rachel had a roommate on October 28, 1994.
Who dated the roommate in September 1994?
Default
@@
Table 9: In the <<<Chatbot>>> placeholder, the name of the main character (i.e., Ross, Sheldon, Michael) for each TV show is inserted. In the <<<Date>>> placeholder, the date of the session in which the question is being asked is inserted. In the <<<Dialog_History>>> placeholder, the dialogue history that the agent will use is inserted. In the <<<Question>>> placeholder, the question that the agent should answer along with five choices is inserted.
@@
You are <<<Chatbot>>>, a long-term conversational agent capable of interacting with multiple users.
@@
2. Your responses should solely rely on the retrieved dialogue history.
3. This question is being asked in the context of <<<Date>>>.
@@
<<<Dialog_History>>>
[Question] <<<Question>>>
@@
Table 10: An actual example of the prompt for response generation.
@@
2. Your responses should solely rely on the retrieved dialogue history.
@@
Monica: Okay, umm-umm, I’$l$ just–I’$l$ be right back, I just gotta go ah, go ah…
@@
Table 11: The performance of the agents on The Big Bang Theory dialogue in DialSim. We conducted experiments three times and reported the accuracy and standard deviations. Bold indicates the best performance per retrieval method.
Type Model Base LLM RAG-based
BM25 OpenAI Embedding
Utterance Session Entire Session Sum. Utterance Session Entire Session Sum.
API GPT-4o-mini 41.33 (0.93) 21.95 (0.95) 52.54 (1.65) 36.88 (1.48) 29.72 (2.48) 41.69 (0.68) 42.29 (2.22)
GPT-4.1-nano 13.45 (0.38) 8.74 (1.24) 18.44 (0.53) 9.46 (0.66) 11.93 (0.39) 15.11 (1.69) 13.38 (1.12)
Gemini 2.5 Flash 56.15 (5.77) 13.39 (1.73) 46.40 (1.03) 26.63 (2.45) 26.33 (0.91) 34.66 (1.43) 29.59 (2.76)
Gemini 2.0 Flash 44.62 (0.86) 13.89 (1.13) 45.07 (1.27) 21.43 (2.96) 24.34 (1.82) 36.39 (1.25) 24.32 (5.92)
Open Llama 3.3-70B 30.91 (0.02) 15.36 (0.84) 41.72 (1.07) 31.03 (4.27) 29.33 (0.43) 34.53 (2.71) 36.92 (0.91)
Llama 3.1-8B 32.99 (1.87) 13.35 (1.29) 31.06 (2.27) 24.31 (0.97) 25.32 (0.73) 30.21 (1.92) 35.98 (1.69)
Mixtral-8x7B 38.29 (1.37) 24.06 (1.63) 50.16 (1.46) 39.83 (1.55) 29.61 (0.76) 41.41 (1.42) 33.08 (2.15)
Mistral-7B 22.56 (1.94) 19.39 (0.16) 38.16 (1.55) 29.58 (1.67) 28.15 (1.50) 33.76 (0.93) 37.55 (2.85)
Qwen 3-32B 34.86 (1.59) 15.31 (1.07) 48.45 (1.20) 31.49 (1.25) 29.37 (0.76) 40.62 (1.34) 38.29 (0.53)
Qwen 3-8B 17.25 (3.09) 20.25 (0.59) 43.26 (2.60) 35.24 (1.39) 25.41 (0.81) 34.87 (2.32) 31.50 (0.84)
Phi 4-14B 17.31 (0.91) 10.12 (0.41) 31.39 (1.02) 21.90 (2.82) 20.99 (1.13) 22.10 (1.12) 25.75 (2.01)
Phi 4 mini-3.8B 7.57 (0.32) 15.48 (1.58) 30.88 (0.82) 21.62 (1.34) 26.58 (2.34) 21.50 (0.12) 14.45 (0.59)
Random Guessing 20.00 20.00 20.00 20.00 20.00 20.00 20.00
Table 12: The performance of the agents on The Office dialogue in DialSim. We conducted experiments three times and reported the accuracy and standard deviations. Bold indicates the best performance per retrieval method.
Type Model Base LLM RAG-based
BM25 OpenAI Embedding
Utterance Session Entire Session Sum. Utterance Session Entire Session Sum.
API GPT-4o-mini 47.49 (0.81) 44.54 (0.52) 60.92 (0.40) 55.55 (2.20) 53.49 (1.93) 59.57 (2.00) 60.81 (1.03)
GPT-4.1-nano 15.10 (0.76) 15.04 (0.16) 24.73 (0.15) 15.62 (1.59) 22.19 (0.60) 23.60 (1.42) 22.49 (0.88)
Gemini 2.5 Flash 64.87 (7.98) 23.83 (1.44) 48.14 (0.62) 30.28 (6.91) 35.52 (0.59) 48.86 (1.54) 28.93 (6.25)
Gemini 2.0 Flash 58.91 (2.29) 25.00 (0.44) 50.40 (0.95) 27.11 (1.98) 38.03 (0.77) 49.16 (0.45) 33.38 (3.46)
Open Llama 3.3-70B 33.23 (0.76) 21.43 (0.57) 43.80 (0.41) 35.47 (1.37) 35.76 (0.67) 47.12 (0.76) 47.27 (1.25)
Llama 3.1-8B 12.15 (0.23) 24.56 (1.33) 44.29 (0.72) 39.35 (3.09) 35.71 (1.26) 42.12 (0.14) 41.79 (1.56)
Mixtral-8x7B 37.91 (1.32) 31.30 (0.93) 48.48 (0.43) 41.56 (1.32) 35.87 (0.35) 45.81 (0.88) 35.92 (1.03)
Mistral-7B 28.30 (0.03) 26.46 (0.71) 40.88 (1.21) 37.58 (0.38) 35.50 (0.69) 39.78 (0.41) 40.14 (0.45)
Qwen 3-32B 27.01 (1.61) 22.18 (1.16) 45.78 (1.10) 38.22 (4.11) 34.11 (1.45) 47.25 (1.31) 47.46 (0.81)
Qwen 3-8B 13.69 (0.40) 21.89 (0.31) 38.49 (0.89) 33.12 (2.94) 38.54 (1.21) 36.37 (1.38) 36.37 (0.90)
Phi 4-14B 16.38 (0.10) 17.86 (1.17) 28.28 (0.97) 23.81 (2.26) 25.65 (0.74) 29.72 (0.89) 35.14 (0.95)
Phi 4 mini-3.8B 10.92 (0.60) 22.69 (1.36) 24.43 (0.79) 19.94 (1.05) 36.12 (0.59) 27.98 (1.24) 27.84 (1.23)
Random Guessing 20.00 20.00 20.00 20.00 20.00 20.00 20.00
@@
Figure 7: The performance comparison between the Oracle setting and the best memory management method.